during the first 4 wks of summer vacation, danny worked at a camp earning $50 per wk. during the remaining 6 wks of vacation, he worked as a stock boy earning $100 per wk. what was his average wkly wage for the summer?
Answer by stanbon(75887)

during the first 4 wks of summer vacation, danny worked at a camp earning $50 per wk.
Total earnings = 4*50 = 200
-------------------------------
during the remaining 6 wks of vacation, he worked as a stock boy earning $100 per wk.
Total earnings = 6*100 = 600
----------------------------------
what was his average wkly wage for the summer?
Equation:
Average = total earned / total weeks = 800/10 = $80.00 per week
=================================================================
Cheers,
Stan H.